A bowl of Halloween candy contains a mixture of Skittles and Kit-Kats. There are 20% more Skittles than Kit-Kats. If there are a total of 198 candies in the bowl, how many are Skittles and how many are Kit-Kats?
Answer by stanbon(75887)

A bowl of Halloween candy contains a mixture of Skittles and Kit-Kats. There are 20% more Skittles than Kit-Kats. If there are a total of 198 candies in the bowl, how many are Skittles and how many are Kit-Kats?
-----------
Equations:
s = 1.2k
s + k = 198
--------
Substitute for "s" and solve for "k":
1.2k + k = 198
2.2k = 198
k = 90 (# of Kit-Kats)
---
s = 1.2*90 = 108 (# of Skittles)
===================
Cheers,
Stan H.
===================